April 2016
Wed
20
Thu
14
Fed Frozen Mysis and Brine 1 tab
Fed Frozen Mysis and Brine 1 tab
Fed Frozen Mysis and Brine 1 tab
Fed Frozen Mysis and Brine 1 tab
Cleaned Glass
Got in tonight and the lobo is really struggling. The flesh is pulling away from the skeleton on one side. I think it might be beyond repair now...
Wed
13
Measured Phosphate at 0.03 ppm
Measured Nitrate at 2 ppm
Measured Calcium at 493 ppm
Measured Salinity at 1.024 SG
Sun
10
Fed Frozen Mysis and Brine 1 tab
Fed Phyto-Plankton 10 ml
Fed Frozen Mysis and Brine 1 tab
Fed Frozen Mysis and Brine 1 tab
Fed Frozen Mysis and Brine 1 tab
Replaced Filter Sock
Cleaned Glass
Fed Phyto-Plankton 10 ml
Fed Feb Live Daphnia 1 tab
Replaced Filter Sock
Changed 25 Liters of water
  15%
Sunday 10th April

Not a great day for the Shaun and Sam aquarium to today. It started with a trip to LFS to collect water for general maintenance and live Daphnia for a feeding treat . When Sam began to do the change she noticed that James Bay (Orange Spotted Goby) had passed away. This was so out of the blue as he'd shown no previous signs of being ill. Sam scooped him out and performed a 25 litre water change, swapped filter sock, blasted rock work and replenished the PhosGuard.

The lobo is also looking its worst. Part of the flesh has now come away from the skeleton, which is frustrating as it had shown signs of recovery earlier in the week and this morning.

Now it's time to look at more Gobies and decide who we're going to replace James Bay with...
Wed
6
Cleaned Glass
Not really required. Since I topped up bio pellets, I've seen a significant improvement with algae on the glass.
Tue
5
Tuesday 5th April

Lobo hasn't got any better up high so moved it back to the sand bed, front central. Not feeling too hopeful of its chances anymore...

Performed full water tests.
Measured pH at 8.2 pH
Measured Temperature at 25.2 C
Measured Alkalinity at 7.0 dKH
Measured Nitrate at 2 ppm
Measured Ammonia at 0.00 ppm
Measured Phosphate at 0.05 ppm
Measured Salinity at 1.024 SG
Measured Calcium at 400 ppm
Really confused by the 2nd Calcium reading. I thought the calcium would have gone up after water change, but it has actually dropped slightly.
Measured Nitrite at 0.00 ppm
Sun
3
Swapped Filter Sock
Cleaned and swapped.
Changed 25 Liters of water
  15%
Sunday 3rd April

Yesterday the lobo had inflated slightly but the hope was short lived as it had receded again this morning.

Battle with GHA is going well, a mixture of manual removal, improved flow, more regular filter sock changes and reduced lighting has seen a much thinner covering.

Went to LFS today and significantly added to CUC. Bought 6 Algae hermits, 2 Cerith snails and 2 Turbo snails so hopefully we'll see further improvements with GHA.

Whilst there we also discussed the lobo and it was suggested that it might not actually be getting enough light. The store runs lights at 100%, I run at 40% so we've decided to move lobo higher in the tank to see if more light improves anything. Time will tell.

Also performed water change and usual maintenance including filter sock swap, glass clean etc.
Sat
2
Added Algae Hermit Crab
To help control Green Hair Algae
Added Cerith Snail
To maintain a clean sand bed
Added Turbo Snail
To help combat Green Hair Algae problem